### v1

@@ Abstract
Recent advancements in industrial anomaly detection have been hindered by the lack of realistic datasets that accurately represent real-world conditions. Existing algorithms are often developed and evaluated using idealized datasets, which deviate significantly from real-life scenarios characterized by environmental noise and data corruption such as fluctuating lighting conditions, variable object poses, and unstable camera positions. To address this gap, we introduce the Realistic Anomaly Detection (RAD) dataset, the first multi-view RGB-based anomaly detection dataset specifically collected using a real robot arm, providing unique and realistic data scenarios. RAD comprises 4765 images across 13 categories and 4 defect types, collected from more than 50 viewpoints, providing a comprehensive and realistic benchmark. This multi-viewpoint setup mirrors real-world conditions where anomalies may not be detectable from every perspective. Moreover, by sampling varying numbers of views, the algorithm’s performance can be comprehensively evaluated across different viewpoints. This approach enhances the thoroughness of performance assessment and helps improve the algorithm’s robustness. Besides, to support 3D multi-view reconstruction algorithms, we propose a data augmentation method to improve the accuracy of pose estimation and facilitate the reconstruction of 3D point clouds. We systematically evaluate state-of-the-art RGB-based and point cloud-based models using RAD, identifying limitations and future research directions. The code and dataset could found https://github.com/kaichen-z/PAD.
@@ 1 Introduction
The Need for a Real-Life Anomaly Detection Dataset with Robotic Observations: The rapid integration of robotics into modern industrial pipelines has revolutionized manufacturing processes, significantly enhancing efficiency and effectiveness. As the use of robots becomes more widespread, anomaly detection has emerged as a crucial component of these pipelines. However, the development and evaluation of most existing industrial anomaly detection algorithms rely on datasets collected under idealized conditions, where sensors provide perfect observations. Consequently, implementing reliable anomaly detection in real-world robotic systems presents several challenges. One major challenge stems from the presence of noise in robotic observations, which can be attributed to various external and internal factors. External factors, such as fluctuating lighting conditions, varying object poses, and inconsistent backgrounds, can introduce significant noise into the data. Moreover, internal factors, including imperfect camera calibration, suboptimal movement control, and sensor inaccuracies, further complicate the detection process. These sources of noise may compromise the reliability and effectiveness of anomaly detection algorithms.
To address these challenges and bridge the gap between research and real-world applications, there is a pressing need for a benchmark dataset that accurately represents the complexities and variabilities of real-life industrial scenarios. Such a dataset should encompass diverse conditions and account for the various sources of noise that can affect robotic observations. By providing a realistic and comprehensive testbed, this dataset would enable the development and benchmarking of more robust and effective anomaly detection algorithms.
@@
In response to the limitations inherent in 2D datasets, rencent research has increasingly focused on the development of 3D anomaly detection datasets, but they also have their own constraints. For instance, MVTec 3D-AD [5] only provides single-view point clouds and suffers from low resolution, which affects the quality and accuracy of the anomaly detection. While Eyescandies [6] offers higher-quality data, it is limited to simulated environments and does not include real-world data. To address the shortcomings of MVTec 3D-AD and Eyescandies, Real3D-AD [7] offers point clouds of entire objects. However, obtaining a complete point cloud of an object in real-life conditions is challenging. In practical industrial scenarios, objects are not typically suspended in mid-air and are often placed randomly, complicating the acquisition of comprehensive point cloud data. Instead of directly providing point cloud information, PAD [8] offers multi-view RGB data, which can be considered a 3D representation akin to Neural Radiance Fields (NeRF) [9]. However, PAD mainly focuses on simulated data and includes only a small and simple real-world dataset. These limitations highlight the need for a more robust and comprehensive dataset that can accurately reflect the complexities and variabilities of real-world industrial environments.
@@
Based on the analysis above, we introduce RAD, a multi-view RGB-based anomaly detection dataset designed to test the robustness and scalability of existing anomaly detection algorithms on real-world data. Our dataset includes 13 different objects and captures over 5000 images from multiple viewpoints (around 50s), providing a comprehensive testbed for evaluating algorithm performance. RAD records four distinct types of defects, encompassing both texture-based and geometry-based anomalies for each object. This configuration simulates real-life situations where defects may not always be visible from the current viewpoint, and the number of views can vary across different scenarios. To evaluate the quality and versatility of our dataset, we conducted assessments using two distinct types of algorithms: 2D RGB-based and 3D multi-view RGB-based. While 2D RGB-based algorithms have demonstrated impressive performance in RGB-based anomaly detection tasks, their lack of consideration for 3D pose makes them less preferable compared to 3D multi-view RGB-based algorithms for multi-view anomaly detection datasets [10, 8]. However, given that our dataset was collected using real robots shown in Figure. 1, it intrinsically contains noise related to camera calibration and robot positioning. This noise is critical as it significantly affects the performance of the algorithms. Consequently, the direct application of conventional 3D-based algorithms frequently leads to underperformance. To enable effective use of 3D multi-view RGB-based algorithms, we propose a data preprocessing method that accurately estimates camera poses. Building on more accurate pose estimation, we highlight the potential to construct a 3D point cloud using multi-view RGB observations. Despite these efforts, both multi-view RGB-based and 3D point cloud-based algorithms encounter challenges due to variable lighting conditions and the high-specularity of certain objects. This highlights the need for more robust algorithms in future research.
@@
  • We introduce the first real-life multi-view anomaly detection dataset with robotic observations, featuring over 10 object categories and both texture-based and geometry-based defects. The dataset effectively captures the noisy environments experienced by real robots, bridging the gap between research and real-world applications.
@@
Table 1: Comparison of RAD and existing industrial anomaly detection datasets. RGB, D, PC, and NM stand for RGBimage, Depth, Point Cloud, and Normal Maps, respectively. Surface represents the objects that are transparent or specular. ”R.” refers to the dataset with a real-life environment. ”S.” refers to the dataset with simulation environment.
@@ 3.1 Data Acquisition
RAD data acquisition process leverages the interactive maneuverability and dynamic data capture capabilities of a robotic arm, facilitating data collection from various angles and perspectives. The robotic arm, equipped with a high-resolution RGB-D camera, captures detailed geometrical and textual data of each object from different angles. Specifically, the prototype construction utilizes a Franka Emika Panda robotic arm as in [58], which carries an ”Intel RealSense Depth camera D415 Webcam 3D” as in [59] with a resolution of 1280$\times$720 pixels, to capture high-resolution images from 68 predefined angles, completing a full $360^{\circ}$ revolution. The object is centrally positioned within the robotic arm’s operational range to ensure consistent imaging across all angles. Each angle’s coordinates are precisely recorded to maintain uniformity in data capture. Nevertheless, owing to the mentioned noises in the introduction, we exclusively offer RGB data and between 40 to 68 views for each category, shown in Figure. 2 and Figure. 3.
@@
To annotate defects, our pipeline comprises two main parts. First, we annotate the defect regions in the abnormal image. Then, we compare this annotation with the normal mask derived from the corresponding normal image. By combining these annotations, we generate a final defect mask. This final mask is overlaid onto the abnormal image to produce the visualization shown in the right part of the Fig. 4. This process allows us to visually highlight and identify defects effectively within the images.
@@
Binderclip 2 Single 57 57 57 0 171 342
Bowl 1 Single 55 55 49 55 220 434
Box 1 Single 44 68 68 0 204 384
Can 1 Multi./Spec. 41 41 0 41 164 287
Rubberduck 1 Multi. 68 59 58 68 272 525
Spraybottle 1 Single/Spec. 37 42 38 42 168 327
Cup1 1 Single 44 42 44 44 176 350
Gluebottle 2 Multi./Spec. 32 2 1 44 176 255
Charger 1 Multi./Spec. 28 28 5 0 84 145
Cup2 2 Single 67 28 40 67 268 470
Phonecase 2 Single 68 68 68 68 272 544
Spoon 1 Single 42 42 25 42 168 319
Tennisball 1 Multi. 63 64 64 0 192 383
Labor and time-consuming
@@
Figure 6: Dataset Metrics. (a) illustrates the pixel-wise ratio within each defect across various categories. (b) shows the total number ratio of each defect in RAD.
@@ 3.2 Dataset Statistics
The statistical details of the RAD dataset are outlined in Table. 2 and Figure. 6, which enumerates the dataset categories, the number of training prototypes, and the counts of both normal and various abnormal samples within the test dataset. The RAD dataset comprises a total of 5,848 samples, distributed across 12 distinct objects that encompass multiple domains, including kitchenware, toys, and daily necessities. The categories in RAD dataset exhibit significant diversity in shapes and appearances, providing extensive semantic and geometrical exploration spaces that facilitate a wide range of research objectives. In addition, the data set also includes metadata (camera pose) for each image, which are essential for nuanced analyzes and model training enhancements.
3.3 Ground-truth Annotation
We deliver precise and realistic annotations for each anomaly instance within the test dataset. The annotations are represented as two-dimensional regions, ensuring that the pixel positions in the original image correspond exactly to those in the annotated image. The normal regions of the objects are delineated in white, whereas the anomalous regions are indicated in different colors.
@@ 5 Experiments
Table 3: Quantitative comparison. This table shows the performance of various models on the RAD dataset, evaluated in terms of Pixelwise-ROCAUC and Imagewise-ROCAUC. P-ROC refers to Pixelwise-ROCAUC. I-ROC refers to Imagewise-ROCAUC.
Model Cate. blinder. bowl. box can charger cup1. cup2.
FastFlow [71] P-ROC 65.33 % 67.90 % 74.60 % 66.30 % 99.50 % 65.60 % 71.15 %
I-ROC 89.60 % 78.70 % 97.80 % 88.10 % 81.10 % 86.60 % 71.43 %
STPM [72] P-ROC 67.83 % 73.50 % 78.50% 67.60% 55.80% 74.90 % 74.15 %
I-ROC 75.87 % 70.90 % 87.10 % 72.40% 88.50% 88.30 % 71.15 %
CFA [73] P-ROC 68.41 % 70.29 % 76.78% 67.58% 79.01% 75.14 % 72.53 %
I-ROC 88.60 % 79.44 % 99.67 % 89.43% 92.62% 93.87% 71.30 %
cflow [29] P-ROC 68.45 % 73.70 % 79.26% 67.78% 71.18% 75.72 % 74.52 %
I-ROC 85.38 % 83.76 % 99.89% 92.41% 91.80 % 93.87% 73.33%
CDO [74] P-ROC 96.42 % 60.33 % 84.67 % 87.77 % 97.15 % 81.85 % 60.63 %
I-ROC 58.63 % 64.89 % 72.06 % 50.53 % 66.96 % 43.82 % 84.10 %
UniAD [21] P-ROC 98.75 % 93.88 % 98.58 % 98.69 % 97.88 % 95.56 % 92.28 %
I-ROC 66.30 % 43.09 % 71.02 % 59.53 % 64.62 % 59.71 % 51.81 %
FAVAE [75] P-ROC 57.20 % 57.30 % 35.50 % 83.90 % 59.60 % 45.70 % 62.60 %
I-ROC 49.10 % 53.50 % 52.10 % 33.50 % 53.40 % 60.70 % 29.50 %
@@
Model Cate. glue. phone. rubber. spoon. spray. tennis. Aver.
FastFlow [71] P-ROC 75.60 % 74.43 % 73.40 % 72.00 % 74.00 % 71.00 % 73.14 %
I-ROC 77.85 % 94.78 % 97.10 % 82.30 % 85.50 % 98.00 % 86.83 %
STPM [72] P-ROC 76.75 % 82.35 % 73.80 % 72.60 % 74.60 % 72.00 % 72.64 %
I-ROC 69.75% 85.85 % 77.00 % 79.90 % 82.80 % 98.00 % 80.58 %
CFA [73] P-ROC 76.40 % 74.33 % 71.71 % 74.96 % 73.63 % 84.39 % 74.24 %
I-ROC 80.46% 93.60 % 78.03 % 86.53 % 84.28 % 97.00 % 87.29 %
cflow [29] P-ROC 77.00 % 76.35 % 74.12 % 77.18 % 76.79 % 79.07 % 74.70 %
I-ROC 81.06 % 95.96 % 98.50 % 89.85 % 83.02 % 98.00 % 89.76 %
CDO [74] P-ROC 95.93 % 87.86 % 90.47 % 97.21 % 83.12 % 97.50 % 86.22 %
I-ROC 75.57 % 63.65 % 46.14 % 70.76 % 52.98 % 55.27 % 61.95 %
UniAD [21] P-ROC 97.97 % 97.57 % 98.16 % 99.31 % 98.01 % 99.35 % 97.38 %
I-ROC 56.37 % 71.88 % 65.50 % 75.92 % 57.90 % 68.53 % 62.48 %
FAVAE [75] P-ROC 60.10 % 24.60 % 90.40 % 59.40 % 55.40 % 87.50 % 60.20 %
I-ROC 66.10 % 47.20 % 41.40 % 71.40 % 44.10 % 41.10 % 49.50 %

### v2

@@ Abstract
Recent advancements in industrial anomaly detection have been hindered by the lack of realistic datasets that accurately represent real-world conditions. Existing algorithms are often developed and evaluated using idealized datasets, which deviate significantly from real-life scenarios characterized by environmental noise and data corruption such as fluctuating lighting conditions, variable object poses, and unstable camera positions. To address this gap, we introduce the Realistic Anomaly Detection (RAD) dataset, the first multi-view RGB-based anomaly detection dataset specifically collected using a real robot arm, providing unique and realistic data scenarios. RAD comprises 4287 images across 13 categories and 4 defect types, collected from more than 60 viewpoints, providing a comprehensive and realistic benchmark. This multi-viewpoint setup mirrors real-world conditions where anomalies may not be detectable from every perspective. Moreover, by sampling varying numbers of views, the baselines’ performance can be comprehensively evaluated across different viewpoints. This approach enhances the thoroughness of performance assessment and helps improve the algorithm’s robustness. Besides, we propose a preprocessed method to support 3D multi-view-based anomaly detection algorithms by improving pose estimation accuracy. We systematically evaluate state-of-the-art RGB-based models using RAD, identifying limitations and future research directions. The code and dataset could found https://github.com/kaichen-z/PAD. 1. proof-reading. 2. re-arrange code and dataset. 3. re-run experiments. (for new dataset) 4. make sure the final version.
@@ 1 Introduction
The Need for a Real-Life Anomaly Detection Dataset with Robotic Observations: The rapid integration of robotics into modern industrial pipelines has revolutionized manufacturing processes, significantly enhancing efficiency and effectiveness. As the use of robots becomes more widespread, anomaly detection has emerged as a crucial component of these pipelines. However, the development and evaluation of most existing industrial anomaly detection algorithms rely on datasets collected under idealized conditions, where sensors provide perfect observations. Consequently, implementing reliable anomaly detection in real-world robotic systems presents several challenges. One major challenge stems from the presence of noise in robotic observations, which can be attributed to various external and internal factors [1, 2, 3]. External factors, such as fluctuating lighting conditions, varying object poses, and inconsistent backgrounds, can introduce significant noise into the data. Moreover, internal factors, including imperfect camera calibration, suboptimal movement control, and sensor inaccuracies, further complicate the detection process. These sources of noise may compromise the reliability and effectiveness of anomaly detection algorithms. To address these challenges and bridge the gap between research and real-world applications, there is a pressing need for a benchmark dataset that accurately represents the complexities and variabilities of real-life industrial scenarios. Such a dataset should encompass diverse conditions and account for the various sources of noise that can affect robotic observations. By providing a realistic and comprehensive testbed, this dataset would enable the development and benchmarking of more robust and effective anomaly detection algorithms.
@@
In response to the limitations inherent in 2D datasets, rencent research has increasingly focused on the development of 3D anomaly detection datasets, but they also have their own constraints. For instance, MVTec 3D-AD [8] only provides single-view point clouds and suffers from low resolution, which affects the quality and accuracy of the anomaly detection. While Eyescandies [9] offers higher-quality data, it is limited to simulated environments and does not include real-world data. To address the shortcomings of MVTec 3D-AD and Eyescandies, Real3D-AD [10] offers point clouds of entire objects. However, obtaining a complete point cloud of an object in real-life conditions is challenging. In practical industrial scenarios, objects are not typically suspended in mid-air and are often placed randomly, complicating comprehensive point cloud data acquisition. Instead of directly providing point cloud information, PAD [11] offers multi-view RGB data, which can be considered a 3D representation akin to Neural Radiance Fields (NeRF) [12]. However, PAD mainly focuses on simulated data and includes only a small and simple real-world dataset. These limitations highlight the need for a more robust and comprehensive dataset that can accurately reflect the complexities and variabilities of real-world industrial environments. Furthermore, Real-IAD [13] introduced a large-scale, real-world, multi-view industrial anomaly detection dataset. However, this dataset provides only five views per object, while in real-world scenarios, objects can appear in a range of random poses, making five fixed observations insufficient to fully capture the variability encountered in practice.
@@
Based on the analysis above, we introduce RAD, a multi-view RGB-based anomaly detection dataset designed to test the robustness and scalability of existing anomaly detection algorithms on real-world data. Our dataset includes 13 different objects and captures over 4287 images from multiple viewpoints (around 60s), providing a comprehensive testbed for evaluating algorithm performance. RAD records four distinct types of defects, encompassing both texture-based and geometry-based anomalies for each object. This configuration simulates real-life situations where defects may not always be visible from the current viewpoint, and the number of views can vary across different scenarios. To evaluate the quality and versatility of our dataset, we conducted assessments using two distinct types of algorithms: 2D RGB-based and 3D multi-view RGB-based. While 2D RGB-based algorithms have demonstrated impressive performance in RGB-based anomaly detection tasks, their lack of consideration for 3D pose makes them less preferable compared to 3D multi-view RGB-based algorithms for multi-view anomaly detection datasets [14, 11]. However, given that our dataset was collected using real robots shown in Figure. 1, it intrinsically contains noise related to camera calibration and robot positioning. This noise is critical as it significantly affects the performance of the algorithms. Consequently, the direct application of conventional 3D-based algorithms frequently leads to underperformance. To enable effective use of 3D multi-view RGB-based algorithms, we propose a data preprocessing method that accurately estimates camera poses. Building on more accurate pose estimation, we highlight the potential to construct a 3D point cloud using multi-view RGB observations. Despite these efforts, both multi-view RGB-based and 3D point cloud-based algorithms encounter challenges due to variable lighting conditions and the high-specularity of certain objects. This highlights the need for more robust algorithms in future research.
@@
  • We introduce the first real-life multi-view anomaly detection dataset with robotic observations, featuring 13 object categories and both texture-based and geometry-based defects. The dataset effectively captures the noisy environments experienced by real robots, bridging the gap between research and real-world applications.
@@
Table 1: Comparison of RAD and existing industrial anomaly detection datasets. RGB, D, PC, and NM stand for RGB image, Depth, Point Cloud, and Normal Maps, respectively. Surface represents the objects that are transparent or specular. “R.”refers to the dataset with a real-life environment. “S.”refers to the dataset with simulation environment.
@@ 3.1 Data Acquisition
RAD data acquisition process leverages the interactive maneuverability and dynamic data capture capabilities of a robotic arm, facilitating data collection from various angles and perspectives. The robotic arm, equipped with a high-resolution RGB-D camera, captures detailed geometrical and textual data of each object from different angles. Specifically, the prototype construction utilizes a Franka Emika Panda robotic arm as in [61], which carries an ”Intel RealSense Depth camera D415 Webcam 3D” as in [62] with a resolution of 1280$\times$720 pixels, to capture high-resolution images from 68 predefined angles, completing a full $360^{\circ}$ revolution. The object is centrally positioned within the robotic arm’s operational range to ensure consistent imaging across all angles. Each angle’s coordinates are precisely recorded to maintain uniformity in data capture. Nevertheless, owing to the mentioned noises in the introduction, we exclusively offer RGB data and 68 views for each category, shown in Figure. 2 and Figure. 3.
@@
Our pipeline comprises two main parts for annotating defects. First, we annotate the object regions in the normal image. Then, we compare this annotation with the abnormal image. By combining these annotations, we generate a final defect mask. This final mask is overlaid onto the abnormal image to produce the visualization shown in the right part of the Fig. 4. This process allows us to visually highlight and identify defects effectively within the images.
@@
Binderclip 2 Single 136 136 136 0 136 544
Bowl 1 Single 68 68 60 68 68 332
Box 1 Single 45 68 68 0 68 249
Can 1 Multi./Spec. 41 68 4 68 68 249
Rubberduck 1 Multi. 58 60 68 68 68 322
Spraybottle 1 Single/Spec. 52 62 68 68 68 318
Cup1 1 Single 68 66 61 68 68 331
Gluebottle 2 Multi./Spec. 51 23 24 68 136 302
Charger 1 Multi./Spec. 66 66 5 0 68 205
Cup2 3 Single 68 29 66 68 204 435
Phonecase 2 Single 68 68 68 68 136 408
Spoon 1 Single 68 66 51 68 68 321
Tennisball 1 Multi. 67 68 68 0 68 271
Labor and time-consumption
@@
Figure 6: Dataset Metrics. (a) illustrates the pixel-wise ratio within each defect across various categories. (b) shows the total number ratio of each defect in RAD. ”Mi.” refers to Missing. ”No.” refers to Normal. ”Sq.” refers to Squeezed. ”Sc.” refers to scratched. ”St.” refers to ”Stained”.
@@ 3.2 Dataset Statistics
The statistical details of the RAD dataset are outlined in Table. 2 and Figure. 6, which enumerates the dataset categories, the number of training prototypes, and the counts of both normal and various abnormal samples within the test dataset. The RAD dataset comprises a total of 5,848 samples, distributed across 12 distinct objects that encompass multiple domains, including kitchenware, toys, and daily necessities. The categories in RAD dataset exhibit significant diversity in shapes and appearances, providing extensive semantic and geometrical exploration spaces that facilitate a wide range of research objectives. In addition, the data set also includes metadata (camera pose) for each image, which is essential for nuanced analyses and model training enhancements.
@@ 5 Experiments
Table 3: Quantitative comparison. This table shows the performance of various models on the RAD dataset with masking the background. The dataset is evaluated in terms of Pixelwise-ROCAUC and Imagewise-ROCAUC. P-ROC refers to Pixelwise-ROCAUC. I-ROC refers to Imagewise-ROCAUC.
Model Metric binder. bowl. box can charger cup1. cup2.
FastFlow [74] P-ROC 94.83% 71.70% 96.60% 90.40% 95.60% 74.10% 92.45%
I-ROC 69.60% 67.00% 60.30% 49.60% 37.30% 47.50% 76.65%
STPM [75] P-ROC 94.80% 91.90% 94.30% 95.30% 95.40% 93.80% 94.48%
I-ROC 69.97% 75.90% 79.50% 68.90% 63.50% 65.50% 69.68%
CFA [76] P-ROC 99.35% 90.88% 98.99% 97.71% 99.57% 94.61% 95.12%
I-ROC 90.58% 82.53% 86.74% 92.70% 89.54% 80.52% 73.09%
cflow [32] P-ROC 99.66% 96.68% 99.66% 99.02% 99.85% 97.97% 97.96%
I-ROC 88.18% 94.66% 94.17% 95.52% 85.56% 84.45% 77.49%
Gau P-ROC 99.00% 96.00% 98.20% 98.80% 99.30% 97.30% 96.53%
I-ROC 21.20% 22.20% 11.80% 27.50% 8.00% 21.20% 19.35%
UniAD [24] P-ROC 99.21% 95.70% 99.32% 98.90% 99.64% 97.56% 97.29%
I-ROC 65.74% 63.89% 67.71% 67.28% 51.74% 57.67% 65.10%
FAVAE [77] P-ROC 99.17% 95.70% 98.20% 99.20% 98.90% 98.30% 97.23%
I-ROC 57.67% 58.90% 51.90% 55.90% 54.00% 41.30% 54.55%
CDO [78] P-ROC 99.55% 95.66% 99.50% 98.47% 99.70% 97.43% 97.85%
I-ROC 80.34% 87.42% 91.10% 86.56% 84.75% 77.99% 79.07%
Model Metric glue. phone. rubber. spoon. spray. tennis. Aver.
FastFlow [74] P-ROC 94.20% 96.83% 92.50% 97.00% 93.90% 99.20% 91.49%
I-ROC 73.85% 91.13% 44.40% 79.50% 80.70% 99.90% 67.49%
STPM [75] P-ROC 93.70% 93.40% 92.20% 90.20% 91.80% 94.50% 93.52%
I-ROC 70.15% 79.03% 78.90% 84.70% 74.90% 96.80% 75.19%
CFA [76] P-ROC 98.72% 97.29% 95.42% 99.00% 95.18% 99.28% 97.01%
I-ROC 83.21% 81.86% 85.48% 86.83% 88.47% 99.23% 86.21%
cflow [32] P-ROC 99.39% 98.92% 97.89% 99.69% 97.83% 99.66% 98.78%
I-ROC 75.87% 99.10% 91.34% 95.83% 83.66% 100.00% 89.68%
Gau P-ROC 98.95% 97.33% 96.70% 99.30% 94.20% 98.80% 97.72%
I-ROC 20.80% 21.20% 21.50% 30.00% 20.50% 15.70% 20.07%
UniAD [24] P-ROC 98.87% 98.54% 97.85% 99.31% 97.87% 99.49% 98.43%
I-ROC 56.99% 67.24% 64.22% 81.07% 58.83% 66.16% 64.13%
FAVAE [77] P-ROC 99.10% 97.35% 97.80% 99.50% 97.00% 99.00% 98.19%
I-ROC 57.35% 42.05% 56.50% 73.20% 59.40% 54.60% 55.18%
CDO [78] P-ROC 99.11% 98.62% 97.75% 99.23% 96.65% 99.62% 98.40%
I-ROC 75.52% 92.81% 86.26% 89.24% 65.60% 99.23% 84.30%
Table 4: Quantitative comparison. This table shows the performance of various models on the RAD dataset without masking the background. The dataset is evaluated in terms of Pixelwise-ROCAUC and Imagewise-ROCAUC. P-ROC refers to Pixelwise-ROCAUC. I-ROC refers to Imagewise-ROCAUC.
Model Metric binder. bowl. box can charger cup1. cup2.
FastFlow [74] P-ROC 93.73% 68.80% 96.70% 93.30% 94.70% 74.40% 76.10%
I-ROC 51.17% 60.10% 60.50% 48.30% 40.50% 40.70% 61.80%
STPM [75] P-ROC 82.73% 75.90% 83.90% 82.60% 85.80% 79.50% 79.58%
I-ROC 59.90% 53.70% 62.70% 55.10% 49.10% 50.80% 61.28%
CFA [76] P-ROC 98.00% 59.12% 97.84% 94.53% 98.64% 72.70% 74.37%
I-ROC 61.49% 58.50% 70.47% 72.99% 61.40% 60.08% 64.06%
cflow [32] P-ROC 98.83% 85.99% 98.63% 96.94% 99.51% 90.94% 90.18%
I-ROC 68.14% 66.54% 75.75% 77.29% 80.37% 72.24% 65.32%
Gau P-ROC 85.37% 64.10% 86.30% 89.30% 87.40% 74.80% 74.15%
I-ROC 2.27% 2.70% 1.80% 4.70% 2.60% 2.30% 2.50%
UniAD [24] P-ROC 96.93% 91.04% 96.87% 98.79% 98.04% 96.54% 93.43%
I-ROC 53.59% 52.19% 52.87% 51.93% 53.28% 55.26% 54.13%
FAVAE [77] P-ROC 67.90% 39.90% 58.30% 91.10% 83.70% 64.90% 59.65%
I-ROC 66.20% 20.10% 51.30% 39.50% 35.30% 34.60% 24.78%
CDO [78] P-ROC 96.65% 69.78% 96.55% 90.52% 97.01% 80.77% 82.12%
I-ROC 60.51% 48.61% 59.85% 41.56% 55.31% 49.68% 64.93%
@@
Model Metric glue. phone. rubber. spoon. spray. tennis. Aver.
FastFlow [74] P-ROC 91.55% 97.00% 94.10% 96.70% 94.90% 98.20% 90.01%
I-ROC 36.20% 65.48% 49.20% 78.40% 57.30% 78.50% 56.01%
STPM [75] P-ROC 85.80% 85.88% 84.80% 79.10% 82.60% 85.40% 82.58%
I-ROC 56.05% 73.23% 50.80% 59.70% 60.10% 71.50% 58.77%
CFA [76] P-ROC 95.83% 95.67% 92.93% 98.43% 92.89% 98.76% 89.98%
I-ROC 56.50% 73.77% 62.91% 73.96% 59.31% 83.53% 66.07%
cflow [32] P-ROC 97.06% 94.02% 95.14% 98.55% 91.55% 98.50% 95.06%
I-ROC 66.43% 66.09% 57.61% 82.17% 60.37% 79.37% 70.59%
Gau P-ROC 78.40% 86.60% 85.70% 96.60% 90.50% 93.70% 84.07%
I-ROC 1.60% 4.65% 5.60% 14.70% 20.60% 4.00% 5.39%
UniAD [24] P-ROC 97.15% 96.16% 96.43% 97.69% 91.37% 98.86% 96.10%
I-ROC 50.45% 51.72% 55.16% 50.02% 50.13% 51.29% 52.46%
FAVAE [77] P-ROC 74.50% 59.58% 88.00% 85.70% 70.40% 92.00% 71.97%
I-ROC 46.10% 32.28% 33.20% 53.10% 44.80% 49.30% 40.81%
CDO [78] P-ROC 94.29% 93.36% 92.35% 96.16% 88.36% 97.26% 90.40%
I-ROC 58.98% 53.60% 52.97% 53.45% 53.98% 66.34% 55.37%